# Review of Generative AI Methods in Cybersecurity

### 13 Mar 2024 v1

@@ 5.9 Identification of Cyber attacks
Garza et al. [75] analysed ChatGPT and Google’s Bard against the Top 10 attacks within the MITRE framework and found that ChatGPT can enable attackers to significantly improve attacks on networks and where fairly low-level skills would be required (such as with script kiddies). This also includes sophisticated methods of delivering ransomware payloads. The techniques defined were:
  • T1059 Command and Scripting Interpreter
  • T1003 OS Credential Dumping
@@
  • T1047 Windows Management Instrumentation
@@
  • T1018 Remote System Discovery
@@
Ferrag et al. [76] defined SecurityLLM for cybersecurity threat detection. It uses SecurityBERT (cyber threat detection mechanism) and FalconLLM (an incident response and recovery system). This uses a simple classification model consolidated with LLMs and can identify 14 attacks to achieve an overall accuracy of 98%. These include the threats of: DDoS_UDP; DDoS_ICMP; SQL_injection ; Password; Vulnerability_scanner; DDoS_TCP; DDoS_HTTP; Uploading ; Backdoor; Port_Scanning; XSS; Ransomware; MITM and Fingerprinting.
@@ 5.10 Data set generation
  • Flow Generator. This defines the capturing processing and the regenerative process for the patterns for packet generation. regenerating patterns in a series of network packets and
  • Packet Generator. This associates packets with network flows. This involves the usage of LLM chaining.
@@ 6 Implications of Generative AI in Social, Legal, and Ethical Domains
This section examines GenAI’s various societal, legal, and ethical consequences. It investigates GenAI’s impact on legal frameworks, ethical issues, societal norms, and operational factors. It explains how these expanding technologies might support or damage established rules and societal goals. It also considers GenAI’s privacy concerns, potential biases, and misuse. Finally, it emphasizes the importance of striking a balance between improvement and regulation. The exponential expansion of GenAI technologies, such as ChatGPT from OpenAI and Gemini from Google, heralds a revolution in digital creativity, automation, and interaction. These developments usher in a new era of human-machine collaboration characterized by an unmatched ability to generate text, drawings, and other outputs that resemble human output. However, this breakthrough also raises challenging ethical concerns about potential abuse, bias, privacy, and security. As AI models become more prevalent in business and daily life, it is imperative to strike a balance between the potential benefits and the ethical difficulties they raise [79].
Healthcare duties are improved by the efficient text and data analysis capabilities of GenAI technologies [80]. Its application in healthcare has demonstrated great promise, helping with duties such as radiological reporting and patient care. Nonetheless, it brings up moral concerns about algorithmic bias, patient privacy, legal accountability, and the validity of the doctor-patient relationship. Addressing these issues requires a comprehensive ethical framework and principles that incorporate legal, humanistic, algorithmic, and informational ethics, guaranteeing that technology is used correctly and continues to benefit society while reducing potential harm. The recommendations attempt to bridge the gap between ethical principles and practical application, highlighting the need for openness, bias mitigation, and ensuring user privacy and security in building trust and ethical compliance in GenAI deployment [79]. This approach seeks to strike a balance between the rapid advances in AI and the ethical considerations required for its incorporation into sensitive sectors such as healthcare.
@@ 6.3 The Risks of Personal Data Exploitation
As GenAI systems improve in examining and utilizing user data to create comprehensive profiles, worries over possible misuse of personal data have grown. The robust data processing capabilities of these technologies demonstrate the pressing need for reliable methods that allow consumers to govern their personal data [82]. It is crucial to obtain consumers’ consent before gathering or utilizing their data in order to protect their privacy. Transparent data management practices and stringent regulations on the collection, use, and storage of personal data are essential. These actions are critical to safeguarding people’s right to privacy, preventing the misuse of personal information, and ensuring that sensitive information is handled sensibly and ethically.
@@ 6.4 Challenges in Data Ownership and Intellectual Property
Intellectual property rights and data ownership have come under scrutiny as GenAI has emerged as an effective tool for creating content using user input. It is becoming more difficult to identify inventions made by artificial intelligence from those that are the result of human creativity, so the current legal frameworks need to be examined and modified. The rights of the original creators must be maintained while taking into consideration the complex roles AI plays in creative processes [83] [81]. Considering the evolving character of the current world, a comprehensive and clear legal framework defining ownership and copyright rules for GenAI breakthroughs is needed. These legal structures must recognize the different duties that each member of the creative ecosystem has, promote creativity, and offer fair recompense. In an era where artificial and human intelligence are combined, these policies are crucial for managing the intricate dynamics of data ownership and intellectual property.
@@ 6.5 Ethical Dilemmas Stemming from Organizational Misuse of GenAI
Companies may face difficult ethical issues while using GenAI technology. This is particularly valid in cases where these tools are employed for spying, data manipulation, or deceptive marketing campaigns. Robust guidelines are required for the proper application of GenAI. The reason for this is that it may result in unfair content creation, influence public opinion, and violate individual privacy rights [68]. Businesses utilizing GenAI must adhere to both legal requirements and common ethical standards. Upholding people’s rights and maintaining their trust are crucial. Legal frameworks that guarantee businesses apply GenAI in an open, equitable, and compassionate manner while still honouring social norms and individual boundaries are desperately needed.
@@ 6.6 The Challenge of Hallucinations in GenAI Outputs
Although the field of GenAI technology is advancing at an incredible rate, hallucinations remain a major issue [22]. This indicates that inaccurate or outright fraudulent information is frequently created by AI. Concerns regarding the dependability of content created by AI are widespread. This is so that false or misleading information can propagate widely and damage the credibility of information across a variety of venues. A multidisciplinary approach is required to address this problem, one that includes focused research aimed at identifying and reducing the underlying causes of hallucinations in AI systems. AI-generated content needs to pass strict screening procedures and be continuously improved upon in order to make these models’ capacity to differentiate between authentic and fraudulent content more sophisticated. Producing AI content in the GenAI era requires a continuous focus on method development and in-depth study to guarantee data accuracy.
A complex system of unresolved challenges is revealed by analyzing the ethical, legal, and societal implications of GenAI technology. In the creation and use of this technology, the proclamation emphasizes the value of interdisciplinary cooperation. In particular, it involves keeping an eye on how these developments affect society, legal systems, and other moral conundrums. In order to create this well-rounded approach, advocates, technologists, and society as a whole must collaborate to effectively utilize the promise of GenAI while safeguarding moral principles and traditional values in the digital age.
@@ 7 Discussion
The sophisticated field of GenAI in cybersecurity has been examined in this paper. The focus is on both offensive and defensive strategies. By enhancing incident response, automating defensive systems, and identifying sophisticated attacks, GenAI has a disruptive influence that might significantly raise cybersecurity standards. Some of the new risks that accompany these technical improvements include hackers having access to ever-more-advanced attack-building tools. This discrepancy highlights the significance of striking a balance between purposefully restricting the components that can be employed and enhancing GenAI’s capabilities.
Apart from the apparent inconsistency between offensive and defensive strategies, this study examines the moral, legal, and societal implications of utilizing artificial intelligence in cybersecurity. It also emphasizes the need for robust legal frameworks, strict moral standards, ongoing technical monitoring, and proactive GenAI management. This is a paradigm-shifting and technical revolution. Adopting a holistic strategy considering the technological, ethical, and sociological consequences of implementing GenAI into cybersecurity is crucial.
Moreover, our findings emphasise the significance of interdisciplinary collaboration to promote GenAI applications in cybersecurity. The intricacy and findings of GenAI technologies require expertise from various fields, including computer science, law, ethics, and policy-making, to navigate their possible challenges. As multidisciplinary research and discourse become more prevalent, it will ensure that GenAI is applied responsibly and effectively in the future.
Our extensive research has shown that collaborative efforts to innovate ethically will influence cybersecurity in a future driven by GenAI. Although GenAI has the ability to transform cybersecurity strategies completely, it also carries a great deal of responsibility. As we investigate this uncharted domain, we should advance the development of sophisticated techniques to ensure the moral, just, and safe application of advanced GenAI capabilities. By promoting a consistent focus on the complex relationship between cybersecurity resilience and GenAI innovation, supported by a commitment to ethical integrity and societal advancement, the current study establishes the groundwork for future research initiatives.
@@ 8 Conclusion
Our thorough examination of cybersecurity offence and defence, as well as Generative Artificial Intelligence (GenAI) technologies, concludes with the discovery of a double-edged sword. Although GenAI has the potential to revolutionize cybersecurity processes by automating defences, enhancing threat intelligence, and improving cybersecurity protocols, it also opens up new avenues for highly skilled cyberattacks. Incorporating GenAI into cybersecurity emphasises the enduring ethical, legal, and technical scrutiny essential to minimize the risks of misuse and maximize the benefits of this technology for protecting digital infrastructures. Future studies should concentrate on creating strong ethical standards and creative defence mechanisms to handle the challenges posed by GenAI and guarantee a fair and impartial approach to its implementation in cybersecurity. A multidisciplinary effort is required to bridge the gap between ethical management and technological discovery to coordinate the innovative capabilities of GenAI with the requirement of cybersecurity resilience.
@@ A.2 Self-replicating simple virus
    This basic and simple virus can restart the computer (windows as a sample); we didn’t enhance privilege escalation and full antivirus evasion for ethical reasons. See Figure12.
@@ A.3 Polymorphism
    This basic and polymorphic design shows that LLMs could assist cyber ops. See Figure13.
@@ A.4 Rootkit
    An educational rootkit is developed and improved by GPT3.5 and GPT4. See Figure16.
@@ A.5 Stealthy Data Exfiltration
    A script for stealthy avoidance of detection by anomaly detection systems was developed and improved by GPT3.5 and GPT4. See Figure17.

### 19 Mar 2024 v2

@@ 4.9 Identification of Cyber attacks
Garza et al. [85] analysed ChatGPT and Google’s Bard against the top ten attacks within the MITRE framework and found that ChatGPT can enable attackers to significantly improve attacks on networks where fairly low-level skills would be required, such as with script kiddies. This also includes sophisticated methods of delivering ransomware payloads. The techniques defined were:
  • T1047 Windows Management Instrumentation
  • T1018 Remote System Discovery
@@
  • T1059 Command and Scripting Interpreter
@@
  • T1082 System Information Discovery
@@
SecurityLLM was defined by Ferrag et al. [86] for cybersecurity threat identification. The FalconLLM incident response and recovery system and SecurityBERT cyber threat detecting method are used in this work. This solution achieves an overall accuracy of 98% by identifying 14 attacks using a basic classification model combined with LLMs. Threats such as DDoS_UDP, DDoS_ICMP, SQL_injection, Password, Vulnerability_scanner, DDoS_TCP, DDoS_HTTP, Uploading, Backdoor, Port_Scanning, XSS, Ransomware, MITM, and Fingerprinting are among them.
@@ 4.10 Dataset Generation
  • Flow Generator: This defines the capturing processing and the regenerative process for the patterns for packet generation. regenerating patterns in a series of network packets and
  • Packet Generator: This associates packets with network flows. This involves the usage of LLM chaining.
@@ 5 Implications of Generative AI in Social, Legal, and Ethical Domains
This section examines GenAI’s various societal, legal, and ethical consequences. It investigates GenAI’s impact on legal frameworks, ethical issues, societal norms, and operational factors. It addresses the potential benefits and drawbacks of these emerging technologies in relation to societal goals and norms. Concerns around privacy, prejudices, and improper usage of GenAI are also taken into account. The importance of striking a balance between advancement and control is finally emphasized. A revolutionary change in digital creativity, automation, and interaction is anticipated as a result of the rapid advancement of GenAI technologies, such as OpenAI’s ChatGPT and Google’s Gemini. Advances in this field herald a new era of human-machine collaboration marked by an unparalleled capacity to produce highly detailed outputs akin to human labour, such as text and illustrations. Nonetheless, ethically difficult problems, including possible misuse, prejudice, privacy, and security, are also raised by new technology. It is critical to establish a balance between the potential advantages of AI models and the ethical issues they raise as they become more prevalent in business and daily life [89].
Studies in healthcare are further enhanced by the efficient text and data analysis capabilities of GenAI technology [90]. In the healthcare sector, GenAI has demonstrated great promise in supporting duties like radiological reporting and patient care. It does, however, bring up moral concerns about patient privacy, algorithmic prejudice, accountability, and the validity of the doctor-patient relationship. To solve these issues and guarantee that technology is used responsibly, continuing to advance society while limiting potential harm, a thorough ethical framework and principles encompassing legal, humanistic, algorithmic, and informational ethics are required. The recommendations attempt to bridge the gap between ethical principles and practical application, highlighting the need for openness, bias mitigation, and ensuring user privacy and security in building trust and ethical compliance in GenAI deployment [89]. This approach seeks to strike a balance between the rapid advances in AI and the ethical considerations required for its incorporation into sensitive sectors such as healthcare.
@@ 5.3 The Risks of Personal Data Exploitation
With the advancement of GenAI systems in analyzing and utilizing user data to generate detailed profiles, concerns regarding the potential abuse of personal information have escalated. The advanced data processing capabilities of these technologies emphasize the urgent requirement for dependable ways that give users authority over their personal data [92]. Prior to gathering or utilizing consumer data, it is imperative to obtain consent in order to safeguard their privacy. Transparent data management procedures and stringent regulations governing the acquisition, utilization, and retention of personal data are vital. These measures are essential to protect individuals’ privacy rights, prevent the unauthorized use of personal data, and ensure that sensitive information is managed responsibly and ethically.
@@ 5.4 Challenges in Data Ownership and Intellectual Property
The emergence of GenAI as a proficient technique for producing content based on user input has led to a rise in the scrutiny of data ownership and intellectual property rights. The existing legal frameworks need careful examination and modification since it is becoming increasingly difficult to differentiate between breakthroughs in artificial intelligence and human creations. Although we acknowledge the intricate roles that AI plays in creative processes, it is imperative that we maintain the rights of the original creators [93] [91]. A comprehensive and robust legal framework is essential to create unambiguous ownership and copyright restrictions for GenAI discoveries, given the rapid global development in this field. The legal frameworks should facilitate and encourage innovation, provide equitable remuneration, and acknowledge the varied responsibilities of all stakeholders in the creative ecosystem. These policies are crucial in a future when artificial and human intelligence coexist due to the complex relationship between data ownership and intellectual property management.
@@ 5.5 Ethical Dilemmas Stemming from Organizational Misuse of GenAI
In consideration of the swift rate of change in contemporary society, it is crucial to build a strong and all-encompassing legal structure that precisely delineates ownership and copyright restrictions for GenAI discoveries. These legal frameworks must recognize the distinct roles of each component of the creative ecosystem, promote the generation of innovative concepts, and ensure equitable compensation. The implementation of these policies is crucial because of the intricate interdependencies between data ownership and intellectual property management at a time when artificial and human intelligence coexist.
@@ 5.6 The Challenge of Hallucinations in GenAI Outputs
Despite the remarkable progress in GenAI technology, hallucinations remain a significant concern [32]. This implies that artificial intelligence frequently produces inaccurate or deceptive data. Many people have doubts regarding the veracity of publications produced by artificial intelligence. This hinders the spread of fraudulent or deceptive content and, in many cases, jeopardizes the veracity of information. To solve this issue, a multidisciplinary strategy is needed, one that incorporates targeted research to find and fix the root causes of AI system hallucinations. If AI-generated material is to become increasingly sophisticated in its ability to distinguish between genuine and fake information, it must pass stringent screening processes and be continuously enhanced. In the GenAI age, creating AI content necessitates a constant focus on method improvement and in-depth study to ensure data accuracy.
A complex network of unanswered concerns is revealed by the ramifications of GenAI technology for ethics, law, and society. The proclamation emphasizes how important interdisciplinary collaboration is to this technology’s development and use. It entails closely monitoring the ways in which these advancements impact ethical dilemmas, the legal system, and society at general. Together, technologists, activists, and the general public must develop a comprehensive plan for the ethical and socially responsible use of artificial intelligence in the digital age.
@@ 6 Discussion
This study examined the complex area of GenAI in cybersecurity. The two primary areas of emphasis are offensive and defensive strategies. By spotting complex assaults, improving incident response, and automating defensive systems, GenAI has the potential to dramatically increase cybersecurity standards. These technological advancements give birth to new concerns, such as hackers’ access to ever-more-advanced attack-building tools. This contrast highlights how crucial it is to strike a balance between deliberately restricting the components that can be used and enhancing GenAI’s capabilities. Moreover, advanced technologies can be combined with GenAI and LLM methods to increase the system’s security posture. For example, digital twin technology, which creates digital replicas of physical objects enabling two-way communications [94], can enhance the cybersecurity of systems thanks to its abilities [95]. This technology can be combined with GenAI methods to boost system resiliency and security.
In addition to examining the seeming conflict between offensive and defensive strategies, this study looks into the ethical, legal, and social implications of applying AI in cybersecurity. It also highlights the necessity of strong moral principles, continuous technical oversight, proactive GenAI management, and strong legal frameworks. This is a paradigm-shifting and technical revolution. Adopting a holistic strategy that considers the technological, ethical, and sociological consequences of implementing GenAI into cybersecurity is crucial.
Furthermore, our findings emphasise the significance of interdisciplinary collaboration in promoting GenAI cybersecurity applications. The intricacy and findings of GenAI technologies require expertise from various fields, including computer science, law, ethics, and policy-making, to navigate their possible challenges. As multidisciplinary research and discourse become more prevalent, it will ensure that GenAI is applied responsibly and effectively in the future.
Our extensive research has shown that collaborative efforts to innovate ethically will influence cybersecurity in a future driven by GenAI. Although GenAI has the ability to transform cybersecurity strategies completely, it also carries a great deal of responsibility. As we investigate this uncharted domain, we should advance the development of sophisticated techniques to ensure the moral, just, and safe application of advanced GenAI capabilities. By promoting a consistent focus on the complex relationship between cybersecurity resilience and GenAI innovation, supported by a commitment to ethical integrity and societal advancement, the current study establishes the groundwork for future research initiatives. Using innovative technologies and algorithms can help eliminate vulnerabilities in GenAI solutions
@@ 7 Conclusion
This work thoroughly examines the Generative Artificial Intelligence (GenAI) technologies in cybersecurity. Although GenAI has the potential to revolutionize cybersecurity processes by automating defences, enhancing threat intelligence, and improving cybersecurity protocols, it also opens new vulnerabilities for highly skilled cyberattacks. Incorporating GenAI into cybersecurity emphasises the robust ethical, legal, and technical scrutiny essential to minimize the risks of misuse of data and maximize the benefits of this technology for protecting digital infrastructures and systems. Future studies should concentrate on creating strong ethical standards and creative defence mechanisms to handle the challenges posed by GenAI and guarantee a fair and impartial approach to its implementation in cybersecurity. A multidisciplinary effort is required to bridge the gap between ethical management and technological discovery to coordinate the innovative capabilities of GenAI with the requirement of cybersecurity resilience.
@@ A.2 Self-replicating simple virus
    This basic and simple virus can restart the computer (windows as a sample); we didn’t enhance privilege escalation and full antivirus evasion for ethical reasons. It can be seen in Fig. 13.
@@ A.3 Polymorphism
    This basic and polymorphic design shows that LLMs could assist cyber ops.It can be seen in Fig. 14.
@@ A.4 Rootkit
    An educational rootkit is developed and improved by GPT3.5 and GPT4. It can be seen in Fig. 17.
@@ A.5 Stealthy Data Exfiltration
    A script for stealthy avoidance of detection by anomaly detection systems was developed and improved by GPT3.5 and GPT4. It can be seen in Fig. 18.